# Understanding Nonlinear Collaboration between Human and AI Agents: A Co-design Framework for Creative Design

### 14 Jan 2024 v1

@@ Abstract.
Nonlinear Human-AI Collaboration, Co-creativity, Creative Design, Creativity Support Tool
copyright: rightsretainedccs: Human-centered computing Collaborative content creationccs: Human-centered computing Empirical studies in HCI
Figure 1. From left to right: (a) pipeline of the traditional linear AI-driven tools; (b) a formative study involving human-human co-design; (c) a nonlinear human-AI co-design framework.
@@ 1. Introduction
Creative design unfolds with an open-ended goal, which cannot be achieved overnight. Despite having an initial idea, designers can continuously generate fresh thoughts when they encounter inspiring stimuli (muller-wienbergen_leaving_2011, ). By fusing the strengths of different thoughts, designers strive to create a unified and cohesive output. With the rapid advancement of AI technology, numerous generative design tools (e.g. Adobe Firefly (adobe, ), Midjourney (midjourney, )) have emerged to aid in the creative design process. These tools can communicate with users and optimize design results according to textual commands. However, they may suffer from poor user experience (e.g. low success rates (buschek_nine_2021, )) due to the ignorance of the natural collaboration process in creative design (howard_describing_2008, ).
@@
According to previous studies (hatchuel_new_2003, ; pugh_total_1991, ; designcouncil, ; gero1990design, ; gero2004situated, ; howard_describing_2008, ), the collaboration workflow in creative activities should be hierarchical and nonlinear. Users are able to progressively refine and align their open-ended requirements as well as remix different design ideas throughout the nonlinear collaboration process (Fig.  1b). In AI-assisted design tools, users often struggle to provide explicit and executable commands, resulting in abstract open-ended requirements (buschek_nine_2021, ). Consequently, the quality of the results generated by AI tools is severely impacted, and it becomes challenging to inspire designers and assist them in remixing designs. Compared to the nonlinear process which involves iterative requirement alignment and option remixing, existing human-AI co-design practices mainly rely on step-by-step command execution, resulting in a more linear workflow (Fig.  1a).
@@
In various related domains, such as requirement analysis (royce1987RA, ; ambreen2018RA, ; maguire2002RA, ; demirel2018RA, ; saraiya2005RA, ) and psychology (2019land, ; flower_cognitive_1981, ; reynante_framework_2021, ), research has focused on studying collaboration patterns among humans and highlighting the importance of requirement alignment and remixing. However, the process of requirement alignment and remixing between humans and AI remains an under-explored problem. To address this problem, we conducted a formative study consisting of two parts. The first part involved pairing participants to collaboratively complete a creative design task in the field of graphic design. The second part involved interviewing the participants to identify pain points in the collaboration process and gather their requirements for AI collaborators. By observing the human-human collaboration process and analyzing the gathered feedback, we summarize five design guidelines for nonlinear AI-assisted tools. These guidelines can be categorized into two main types: implementing nonlinear characteristics in human-AI collaboration (e.g., providing alternative design solutions for inspiration) and improving the shortcomings of human-human collaboration (e.g., achieving real-time visual outcomes).
@@
Based on these guidelines, we propose a novel framework (Fig.  1c) that introduces the nonlinear human-AI collaboration for creative design tasks. Our framework exhibits three important features: providing different options for each command, explaining how each option is derived from the user’s command, and allowing the remixing of any existing outputs. To validate the effectiveness of the framework, we develop a proof-of-concept prototype (POC), OptiMuse, and evaluate it through a comparative experiment using the Wizard-of-Oz method. To establish a baseline tool, we adopt the state-of-the-art tool, Office 365 Copilot for PowerPoint (copilot, ), which follows the linear human-AI workflow. Copilot quickly provides users with an optimal answer to their commands without explaining how the answer is inferred from the user’s instructions. We recruited 12 participants to use OptiMuse and Copilot to complete creative design tasks. The experiment results indicate that OptiMuse significantly improves users’ task completion rates. Upon analyzing the interview data, we identify four distinct user expectations regarding the roles of AI agents.
Our main contributions are as follows:
@@
  • We conducted a formative study to investigate the co-design process between human collaborators and formulated a set of guidelines for establishing human-AI co-design tools.
@@
  • We conducted a comparative study to evaluate our framework and demonstrate the usage of OptiMuse.
@@ 2. Related Work
According to the well-established taxonomy (howard_describing_2008, ; wang_literature_2017, ), we introduce relevant studies with regards to (i) the concept and practice of the creative design process, (ii) the methodologies adopted by HCI researchers in developing creativity-support tools, and (iii) the strategies utilized in human-human collaboration to foster creativity.
@@ 2.1. Creative Design Process
In the design process, creativity plays an irreplaceable role in promoting innovation (mumford_creativity_1988, ; yusuf2009creativity, ). The investigation of the creative process, as opposed to the conventional design approach, piques people’s interest. Some models (hatchuel_new_2003, ; pugh_total_1991, ; designcouncil, ; gero1990design, ; gero2004situated, ; howard_describing_2008, ) incorporate nonlinear design workflow that might give rise to creativity. For instance, the “C-K theory” (hatchuel_new_2003, ) portrays design as a nonlinear navigation within two spaces: the concept space (C) and the knowledge space (K). The “concept space” refers to a space of basic semantic meanings in the design process while the “knowledge space” means a logical space of designers’ “propositions”. As this model depicts, the design process is dynamic, evolving between the two spaces. Another design model is the Double Diamond (designcouncil, ) which, as a typical divergent-convergent model, includes exploring widely (divergent thinking) and taking focused action (convergent thinking). Moreover, the looped arrays in the model remind the designers that practices toward innovation are featured by repeated evaluation and iteration. Gero’s function, behavior, and structure (FBS) model (gero1990design, ) comprises eight fundamental design processes, with particular emphasis on the three reformulation processes. These processes underscore a dynamic design approach by emphasizing the significance of making changes in response to identified dissatisfaction. Furthermore, the situated FBS framework (gero2004situated, ) sets a new foundation for intelligent design systems by leveraging the “reformulation” of agents in an open, dynamic world. Combining the view of cognitive psychology, T. J. Howard et al. (howard_describing_2008, ) mapped analysis, generation, and evaluation as the consensus of the creative process into the FBS model. The recognition and emphasis on the nonlinear nature of the creative process have been evident in various design models. However, the advancement of AI has rendered the need for a “renaissance” in a creative design framework with a focus on human-AI collaboration.
Certainly, while these nonlinear models may capture the essence of the creative design process, their high-level descriptions may not seamlessly translate to the intricacies of real-world design (howard_describing_2008, ). Rather than strictly adhering to one of these nonlinear creative design models, we aim to glean insights directly from design practices, enabling the formulation of models that better align with real-world scenarios.
@@ 2.2. From Creativity Support Tool to Human-Computer Co-creativity
How can computers actively engage in the creative design process to ignite and enhance creativity? The HCI community has investigated this issue based on the design lessons about developing creativity-support tools (wang_literature_2017, ; Althuizen_Creative_2016, ; clark_creative_2018, ; kulkarni2023word, ; oh_i_2018, ). Tools that support creative tasks have been widely studied (wang_literature_2017, ; muller-wienbergen_leaving_2011, ), including approaches supporting the creative process of finding problems, ideas, and solutions (wang_literature_2017, ). For example, based on the idea of “machine-in-the-loop” Clark et al.’s machine agent (clark_creative_2018, ) provides writing suggestions for human users to choose from, and users are always in control of the final results. Oh et al. (oh_i_2018, ) developed DuetDraw, a creative drawing tool equipped with AI, to provide drawing ideas and directly draw on the canvas. Moreover, by evaluating a set of AI-steering music design tools, Louie et al. (louie_novice-ai_2020, ) reported positive user feedback on human-AI collaboration. Kulkarni et al. (kulkarni2023word, ) demonstrated that generative text-to-image models act as a helpful solution for novice users in design exploration and can promote human-AI collaboration. Although design theories acknowledge the non-linear navigation between design stages (hatchuel_new_2003, ; pugh_total_1991, ; designcouncil, ; gero1990design, ; gero2004situated, ; howard_describing_2008, ), these creativity-support tools often concentrate on examining one or two stages individually (wang_literature_2017, ). This limitation impedes the transformation of computers from being viewed as mere tools for creative support to being perceived as inherently creative entities. Recently, AI agents have become more extensively integrated into the creative process, which requires new studies to promote nonlinear human-AI collaboration.
@@
Various HCI theories (lubart2005can, ; wang_literature_2017, ; muller-wienbergen_leaving_2011, ; negrete2014apprentice, ; kantosalo2016modes, ; guzdial2019interaction, ; rezwana2022understanding, ) have been conducted to provide guidlines for human-AI co-creativity. The apprentice framework (negrete2014apprentice, ) analyzes creative processes as a collaborative effort between humans and computers, assigning them distinct roles and tasks. It identifies four creativity pieces emerging during the creative design process and five roles a computer can play to produce them. From a computational perspective, Kantosalo et al. (kantosalo2016modes, ) defined alternating co-creativity where human and AI equally take turns to finish creative tasks. The process lifts strict requirements to complete creative agent who must be capable of identification, generation, and evaluation. Moreover, Guzdial et al. (guzdial2019interaction, ) presented a general framework for turn-based human-AI co-creativity by identifying components and design reflections. A user study conducted by Rezwana et al. (rezwana2022understanding, ) demonstrates that participants perceive AI as more human-like—depicting qualities when facilitating AI-to-human communication. These works mainly conceive of the nonlinear creative design process as a straightforward, turn-based interaction between users and computers. By thoroughly examining human-human creativity practices, with a particular focus on nonlinearity, our study introduces a novel human-AI co-design framework that more accurately reflects the nonlinear navigation through the various stages of creative design.
@@ 2.3. Collective Creativity
As Maher (maher2012computational, ) stated, “Creativity can ascribed to a computational agent, an individual person, collectives of people and agents and/or their interaction.” The social nature of creativity, which highlights that much of our creative output is a result of interaction and communication with others (fischer2004social, ), emphasizes the significance of collective creativity as a method that promotes innovation through collaborative efforts (Inakage2007collective, ). For example, the TickRok music Duets enable a collective exploration of design space to increase ideation creativity (O'Toole2023collaborative, ). Specifically, the “cumulative” evolution of creativity is achieved in the parallel interaction of multiple design versions. The micro-task approach (Girotto2016collective, ), another collective creativity strategy, enables synergistic collaboration by defining sub-tasks and assigning them to skilled and motivated communities. In real practice, group collaboration is accompanied by “remixing”, which is important in the creative processes (Dynamicland, ; scratch, ; bruckman_community_1998, ). Many design tools allow designers to merge their human collaborators’ ideas into their creative output iteratively (dasgupta_remixing_2016, ; bruckman_community_1998, ; marlow_rookie_2014, ). There has been a surge in research interest in exploring methods to foster innovative remixing (tan2022conflict, ; Cheliotis2014remix, ; Oehlberg2015remix, ; yue2019improving, ). For instance, through an empirical study on the Thingiverse (a 3D printing community), Tan et al. (tan2022conflict, ) revealed the contradictory impact of knowledge endowment heterogeneity on users’ quality and quantity of remixing. However, there is a gap in understanding how AI can support people in remixing to enhance creativity in human-AI collaboration. Creativity is a crucial aspect of graphic design, enabling designs to stand out from their competitors and capture viewers’ attention (graphicdesign, ). Our study aims to enhance the understanding of collective creativity by proposing a collaboration framework for human-AI creative design and investigating how individual designers re-mix alternative designs when working on AI-assisted graphic designs.
3. Study I: Understand the Nonlinear Process in Human-Human Creative Design
@@
The aim of Study I is to establish guidelines for fostering better collaboration between human and AI agents during the creative design process. Building upon the insights from existing study (wang_human-human_2020, ) that emphasize the integration of human-centered design philosophy into complex human activities, we first sought to gain a deeper understanding of human-human collaboration in creative design (Fig. 2). We observe users’ co-design process when they face-to-face optimized graphic design. Our expected results mainly focus on two aspects: (1) How people express their vague requirements and conclude specific operations related to the requirements in natural language? (2) How these vague requirements transform to specific operations?
@@ 3.1. Preconsideration of Experiment Settings
The creative process is a purpose-driven journey in which individuals construct a hierarchical network of goals to serve as their guiding framework(flower_cognitive_1981, ). Requirement analysis (RA), namely a thorough analysis of the goals, is important during the process because it helps ensure that the desired objectives and outcomes of a project are clearly understood and defined. While RA in the context of design modification has not been extensively studied, it has been a topic of longstanding discussion in other domains, particularly in software development (ambreen2018RA, ; demirel2018RA, ; maguire2002RA, ; royce1987RA, ; saraiya2005RA, ). Our formative study design incorporates the knowledge derived from RA.
To explore the possible task division between humans and AI agents, we followed a formative study strategy presented by Pan et al.’s study (pan_human-computer_2023, ). We simultaneously observed the behaviors of designers and design assistants. Also, the role-playing strategy helped participants to recall relevant experiences. To obtain guidance from the collaborative process, it is important to develop scenarios that are conducive to collaboration. We employed a role-playing strategy and set up the best collaboration scenario during the formative study. A pilot study was conducted among three experienced design experts to help us determine the procedure. They had professional experience in design for 3, 4, and 5 years, respectively. Two experts and the experimenter gathered in person, assembling around a laptop. They shared their screens with another expert who participated remotely through an online meeting platform. The experts engaged in discussions on design enhancements collectively without a predefined procedure.
Based on the natural co-design process shown in the pilot study, we established favorable collaborative conditions for the participants. Due to the limitations of online meetings, which hindered effective communication in co-design, the modification sessions were conducted in person, enabling multimodal interactions and real-time communication. Additionally, we allocated time for content comprehension to ensure a fundamental understanding of the target design. Afterward, designers brainstormed individually in preparation for the discussion.

### 20 Mar 2024 v3

@@ Abstract.
Human-AI Co-creativity, Creative Design, Creativity Support Tool
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642812isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Collaborative content creationccs: Human-centered computing Empirical studies in HCI
Figure 1. (a) pipeline of the traditional linear AI-driven tools; (b) a formative study involving human-human co-design; (c) a nonlinear human-AI co-design framework.
@@ 1. Introduction
Creative design unfolds with an open-ended goal, which cannot be achieved overnight. Despite having an initial idea, designers can continuously generate fresh thoughts when they encounter inspiring stimuli (Müller-Wienbergen et al., 2011). By fusing the strengths of different thoughts, designers strive to create a unified and cohesive output. With the rapid advancement of AI technology, numerous generative design tools (e.g. Adobe Firefly (Adobe, 2024), Midjourney (Midjourney, 2022)) have emerged to aid in the creative design process. These tools can communicate with users and optimize design results according to textual commands. However, they may suffer from poor user experience (e.g. low success rates (Buschek et al., 2021)) due to the ignorance of natural collaboration process in creative design (Howard et al., 2008).
@@
According to previous studies (Hatchuel and Weil, 2003; O’Connor, 1991; Council, 2024; Gero, 1990; Gero and Kannengiesser, 2004; Howard et al., 2008), the collaboration workflow in creative activities should be nonlinear. Users progressively refine and align their open-ended requirements as well as remix different design ideas throughout the nonlinear collaboration process (Fig.  1b). When using AI-assisted design tools, users often struggle to provide explicit and executable commands, resulting in abstract open-ended requirements (Buschek et al., 2021). Consequently, the quality of the results generated by AI is severely impacted, and it becomes challenging to inspire designers and assist them in remixing design ideas. Compared to the nonlinear process which involves requirement alignment and option remixing, existing human-AI co-design practices mainly rely on command execution, resulting in a more linear workflow (Fig.  1a).
@@
In various related domains, such as requirement analysis (Royce, 1987; Ambreen et al., 2018; Maguire and Bevan, 2002; Demirel and Das, 2018; Saraiya et al., 2005) and psychology (Frich et al., 2019; Flower and Hayes, 1981; Reynante et al., 2021), research has focused on studying collaboration patterns among humans and highlighting the importance of requirement alignment and remixing. However, the process of requirement alignment and remixing between humans and AI remains an under-explored problem. To address this problem, we conducted a formative study consisting of two parts to learn from the human-human collaboration. Taking graphic design as a representative scenario, the first part involved pairing participants to collaboratively complete a creative design task. The second part involved interviewing the participants to identify challenges in the collaboration process and gather their requirements for AI collaborators. By observing the human-human collaboration process and analyzing the gathered feedback, we summarize five design requirements for nonlinear AI-assisted tools. These design requirements can be categorized into two main types: implementing nonlinear characteristics in human-AI collaboration (e.g., providing alternative design solutions for inspiration) and improving the shortcomings of human-human collaboration (e.g., achieving real-time visual outcomes).
@@
Based on the formative study, we propose a novel framework (Fig.  1c) that introduces the nonlinear human-AI collaboration for creative design tasks. Our framework incorporates ten specific actions for AI agents centered around five areas (flexible communication, clarification, alternative solutions, operational strategies, and visual outcomes) corresponding to the design requirements. To validate the effectiveness of the framework, we develop a proof-of-concept prototype (POC), OptiMuse, and evaluate it through a comparative experiment using the Wizard-of-Oz method. To establish a baseline tool, we adopt the state-of-the-art tool, Office 365 Copilot for PowerPoint (Microsoft, 2023), which follows the linear human-AI workflow. Copilot provides users with an optimal output to their commands without explaining how it is inferred from the user’s instructions. We recruited 12 participants to complete creative design tasks using OptiMuse and Copilot. The experiment results indicate that OptiMuse significantly improves users’ task completion rates. Upon analyzing the interview data, we identify four expected roles of AI agents. Our main contributions are as follows:
@@
  • We conducted a formative study to investigate the co-design process between human collaborators and formulated a set of design requirements for establishing human-AI co-design tools.
@@
  • We conducted a comparative study to validate our framework and evaluate the usability of OptiMuse.
@@ 2. Related Work
According to the well-established taxonomy (Howard et al., 2008; Wang and Nickerson, 2017), we introduce relevant studies with regards to (1) the concept and practice of the creative design process, (2) the methodologies adopted by HCI researchers in developing creativity-support tools, and (3) the strategies utilized in human-human collaboration to foster creativity.
@@ 2.1. Creative Design Process
In the design process, creativity plays an irreplaceable role in promoting innovation (Mumford and Gustafson, 1988; Yusuf, 2009). The investigation of the creative process, as opposed to the conventional design approach, piques people’s interest. Some models (Hatchuel and Weil, 2003; O’Connor, 1991; Council, 2024; Gero, 1990; Gero and Kannengiesser, 2004; Howard et al., 2008) incorporate nonlinear design workflow that might give rise to creativity. For instance, the “C-K theory” (Hatchuel and Weil, 2003) portrays design as a nonlinear navigation within two spaces: the concept space (C) and the knowledge space (K). The “concept space” refers to a space of basic semantic meanings in the design process while the “knowledge space” means a logical space of designers’ “propositions”. As this model depicts, the design process is dynamic, evolving between the two spaces. Another design model is the Double Diamond (Council, 2024) which, as a typical divergent-convergent model, includes exploring widely (divergent thinking) and taking focused action (convergent thinking). Moreover, the looped arrays in the model remind the designers that practices toward innovation are featured by repeated evaluation and iteration. Gero’s function, behavior, and structure (FBS) model (Gero, 1990) comprises eight fundamental design processes, with particular emphasis on the three reformulation processes. These processes underscore a dynamic design approach by emphasizing the significance of making changes in response to identified dissatisfaction. Furthermore, the situated FBS framework (Gero and Kannengiesser, 2004) sets a new foundation for intelligent design systems by leveraging the “reformulation” of agents in an open, dynamic world. Combining the view of cognitive psychology, T. J. Howard et al. (Howard et al., 2008) mapped analysis, generation, and evaluation as the consensus of the creative process into the FBS model. The recognition and emphasis on the nonlinear nature of the creative process have been evident in various design models. However, the advancement of AI has rendered the need for a “renaissance” in a creative design framework with a focus on human-AI collaboration.
Certainly, while these nonlinear models may capture the essence of the creative design process, their high-level descriptions can not seamlessly translate to the intricacies of real-world design (Howard et al., 2008), especially for human-AI collaboration practices. Therefore, rather than strictly adhering to one of these nonlinear creative design models, we aim to glean insights directly from human-human design practices among design experts, enabling the formulation of models that better align with real-world scenarios.
@@ 2.2. From Creativity Support Tool to Human-Computer Co-creativity
How can computers actively engage in the creative design process to ignite and enhance creativity? The HCI community has investigated this issue based on the design lessons about developing creativity-support tools (Wang and Nickerson, 2017; Althuizen and Reichel, 2016; Clark et al., 2018; Kulkarni et al., 2023; Oh et al., 2018). Tools that support creative tasks have been widely studied (Wang and Nickerson, 2017; Müller-Wienbergen et al., 2011), including approaches supporting the creative process of finding problems, ideas, and solutions (Wang and Nickerson, 2017). For example, based on the idea of “machine-in-the-loop” Clark et al.’s machine agent (Clark et al., 2018) provides writing suggestions for human users to choose from, and users are always in control of the final results. Oh et al. (Oh et al., 2018) developed DuetDraw, a creative drawing tool equipped with AI, to provide drawing ideas and directly draw on the canvas. Moreover, by evaluating a set of AI-steering music design tools, Louie et al. (Louie et al., 2020) reported positive user feedback on human-AI collaboration. Kulkarni et al. (Kulkarni et al., 2023) demonstrated that generative text-to-image models act as a helpful solution for novice users in design exploration and can promote human-AI collaboration. Although design theories acknowledge the non-linear navigation between design stages (Hatchuel and Weil, 2003; O’Connor, 1991; Council, 2024; Gero, 1990; Gero and Kannengiesser, 2004; Howard et al., 2008), these creativity-support tools often concentrate on examining one or two stages individually (Wang and Nickerson, 2017) without a recognition of the nonlinear navigation between these stages. This limitation impedes the transformation of computers from being viewed as mere tools for creative support to being perceived as inherently creative entities.
@@
Recently, AI agents have become more integrated into the creative process, requiring new studies to promote natural human-AI collaboration. Various HCI theories (Lubart, 2005; Wang and Nickerson, 2017; Müller-Wienbergen et al., 2011; NEGRETE YANKELEVICH and MORALES ZARAGOZA, 2014; Kantosalo and Toivonen, 2016; Guzdial and Riedl, 2019; Rezwana and Maher, 2022) have been conducted to provide guidelines for human-AI co-creativity. The apprentice framework (NEGRETE YANKELEVICH and MORALES ZARAGOZA, 2014) analyzes creative processes as a collaborative effort between humans and computers, assigning them distinct roles and tasks. It identifies four creativity pieces emerging during the creative design process and five roles a computer can play to produce them. From a computational perspective, Kantosalo et al. (Kantosalo and Toivonen, 2016) defined alternating co-creativity where human and AI equally take turns to finish creative tasks. The process lifts strict requirements to complete creative agent who must be capable of identification, generation, and evaluation. Moreover, Guzdial et al. (Guzdial and Riedl, 2019) presented a general framework for turn-based human-AI co-creativity by identifying components and design reflections. A user study conducted by Rezwana et al. (Rezwana and Maher, 2022) demonstrates that participants perceive AI as more human-like—depicting qualities when facilitating AI-to-human communication. However, these works mainly conceive of the nonlinear creative design process as a straightforward, turn-based interaction between users and computers. By thoroughly examining human-human creativity practices, with a particular focus on nonlinearity, our study introduces a novel human-AI co-design framework that more accurately reflects the nonlinear navigation through the various stages of creative design.
@@ 2.3. Collective Creativity
As Maher (Maher, 2012) stated, “Creativity can ascribed to a computational agent, an individual person, collectives of people and agents and/or their interaction.” The social nature of creativity, which highlights that much of our creative output is a result of interaction and communication with others (Fischer, 2004), emphasizes the significance of collective creativity as a method that promotes innovation through collaborative efforts (Inakage, 2007). For example, the TickRok music Duets enable a collective exploration of design space to increase ideation creativity (O’Toole, 2023). Specifically, the “cumulative” evolution of creativity is achieved in the parallel interaction of multiple design versions. Another example is the micro-task approach (Girotto, 2016), a collective creativity strategy that enables synergistic collaboration by defining sub-tasks and assigning them to skilled and motivated communities. Group collaboration that sparks creativity is always accompanied by “remixing”, which is important and frequently observed in real-world creativity practices (Group, nd; Foundation, 2022; Bruckman, 1998). Many design tools allow designers to merge their human collaborators’ ideas into their creative output iteratively (Dasgupta et al., 2016; Bruckman, 1998; Marlow and Dabbish, 2014). Also, there has been a surge in research interest in exploring methods to foster innovative remixing (Tan et al., 2022; Cheliotis et al., 2014; Oehlberg et al., 2015; Yue et al., 2019). For instance, through an empirical study on the Thingiverse (a 3D printing community), Tan et al. (Tan et al., 2022) revealed the contradictory impact of knowledge endowment heterogeneity on users’ quality and quantity of remixing. However, there is a gap in understanding how AI can support people in remixing to enhance creativity in human-AI collaboration. To fill this gap, we started with a typical scenario of creative design. Creativity is a crucial aspect of graphic design, enabling designs to stand out from their competitors and capture viewers’ attention (Lopes et al., 2022). Our study aims to enhance the understanding of collective creativity by proposing a collaboration framework for human-AI creative design and investigating how individual designers re-mix alternative designs when working on AI-assisted graphic designs.
3. Study I: Understand the Nonlinear Process in Creative Design
@@
The aim of Study I is to establish guidelines for fostering better collaboration between human and AI agents during the creative design process. Building upon the insights from existing study (Wang et al., 2020) that emphasize the integration of human-centered design philosophy into complex human activities, we first sought to gain a deeper understanding of human-human collaboration in creative design  (Fig. 2). We observe users’ co-design process when they face-to-face optimize graphic design. Our expected results focus on:
  1. (1)
    How do designers express vague requirements and the related specific operations in natural language?
  2. (2)
    How do vague requirements transform into specific operations?
@@ 3.1. Preconsideration of Experiment Settings
The creative process is a purpose-driven journey in which individuals construct a hierarchical network of goals to serve as their guiding framework(Flower and Hayes, 1981). Requirement analysis (RA), a thorough analysis of the goals, is important during the process because it helps ensure that the desired objectives and outcomes of a project are clearly understood and defined. While RA in the design process has not been extensively studied, it has been a topic of longstanding discussion in other domains, particularly in software development (Ambreen et al., 2018; Demirel and Das, 2018; Maguire and Bevan, 2002; Royce, 1987; Saraiya et al., 2005). Our formative study design incorporates the knowledge derived from RA.
To explore the possible task division between human and AI agent, we followed a formative study strategy presented by Pan et al.’s study (Pan et al., 2023). We simultaneously observed the behaviors of designers and design assistants. To obtain guidance from the collaborative process, it is important to develop scenarios that are conducive to collaboration. We employed a role-playing strategy and set up the best collaboration scenario during the formative study. The role-playing strategy also helped participants to recall relevant experiences.
A pilot study was conducted among three experienced design experts to help us determine the details of the procedure. They had professional experience in design for 3, 4, and 5 years, respectively. Two experts and the experimenter gathered in person, assembling around a laptop. They shared their screens with another expert who participated remotely through an online meeting platform. The experts engaged in discussions on design enhancements collectively without a predefined procedure. Based on the natural co-design process shown in the pilot study, we established favorable collaborative conditions for the participants. Because online meetings would hinder effective communication in co-design, the modification sessions were conducted in person, enabling multimodal interactions and real-time communication. Additionally, we allocated time for content comprehension to ensure a fundamental understanding of the target design. Afterward, designers brainstormed individually in preparation for the discussion.